OpenStudy (anonymous):

PLZ Help! will medal for answer.... find fog, gof, fof and gog and domain f(x)=sqare root x g(x)=2/(x-4)

OpenStudy (anonymous):

someone plz help!!!

HanAkoSolo (jamierox4ev3r):

I'm sorry I'm not quite sure how to do this stuff.. sorry that the people in math chat were not being incredibly nice, my sincerest apologies

OpenStudy (anonymous):

not ur fault I thought this group was to help but Its just stupid no1 even bothers to help

HanAkoSolo (jamierox4ev3r):

people will help…think of the groups as a bit of a study break :) to get some help, tag people with a high smartscore of 90 and over…. and I recommend you bumping your question so more people can see it

HanAkoSolo (jamierox4ev3r):

and once again, I apologize for not being to help you…not quite sure how to do this

OpenStudy (anonymous):

thanks @Jamierox4ev3r

OpenStudy (ymendoza88):

f o g= square root (2/x-4), g o f= 2/((square root x)-4), f o f= square root(square root x), g o g= 2/((2/x-4)-4) and then solve for them

OpenStudy (ranga):

\[f(x) = \sqrt{x} \quad Domain \quad x \ge 0\]\[g(x) = \frac{ 2 }{ x-4 } \quad Domain \quad x \neq 4\]

OpenStudy (ranga):

\[(fog)(x) = f(g(x)) = \sqrt{\frac{ 2 }{ x-4 }}~~Domain~~x \gt 4\]

OpenStudy (ranga):

\[(gof)(x) = g(f(x)) = {\frac{ 2 }{ \sqrt{x}-4 }}~~Domain~~x \ge 0, x \ne 16\]

OpenStudy (ranga):

Complete the rest in a similar manner. If they want the domain in interval notation, then write x > 4 as (4, infinity) and write x >= 0, x not equal to 16 as: [0, 16) U (16, infinity) where U is the union symbol.

OpenStudy (anonymous):

thank u so much @ranga